### v1

# MSL: Not All Tokens Are What You Need for Tuning LLM as a Recommender

Abstract.
Large language models (LLMs), known for their comprehension capabilities and extensive knowledge, have been increasingly applied to recommendation systems (RS). Given the fundamental gap between the mechanism of LLMs and the requirement of RS, researchers have focused on fine-tuning LLMs with recommendation-specific data to enhance their performance. Language Modeling Loss (LML), originally designed for language generation tasks, is commonly adopted. However, we identify two critical limitations of LML: 1) it exhibits significant divergence from the recommendation objective; 2) it erroneously treats all fictitious item descriptions as negative samples, introducing misleading training signals.
To address these limitations, we propose a novel Masked Softmax Loss (MSL) tailored for fine-tuning LLMs on recommendation. MSL improves LML by identifying and masking invalid tokens that could lead to fictitious item descriptions during loss computation. This strategy can effectively avoid the interference from erroneous negative signals and ensure well alignment with the recommendation objective supported by theoretical guarantees. During implementation, we identify a potential challenge related to gradient vanishing of MSL. To overcome this, we further introduce the temperature coefficient and propose an Adaptive Temperature Strategy (ATS) that adaptively adjusts the temperature without requiring extensive hyperparameter tuning. Extensive experiments conducted on four public datasets further validate the effectiveness of MSL, achieving an average improvement of 42.24% in NDCG@10. The code is available at https://github.com/WANGBohaO-jpg/MSL.
Sequential Recommendation, Large Language Model, Loss Function
ccs: Information systems Recommender systems
1. Introduction
Large Language Models (LLMs) have showcased exceptional capabilities in content comprehension and leveraging extensive knowledge, thereby catalyzing a revolution in artificial intelligence (Achiam et al., 2023). Recently, LLMs have been extensively applied in the field of Recommender Systems (RS) (Wu et al., 2024). A prominent strategy involves directly leveraging LLMs as recommenders — organizing users’ historical interactions as language prompts and instructing LLMs to deduce users’ preferences for predicting future interactions (Hou et al., 2024; Wang and Lim, 2023; Gao et al., 2023b; Wang et al., 2024d; Liu et al., 2024a). This paradigm has demonstrated enhanced few-shot ability (Hou et al., 2024; Wang and Lim, 2023), generalization (Kolb, 2024), explainability (Gao et al., 2023b), and impressive recommendation performance (Bao et al., 2023b).
Figure 1. Language modeling loss can erroneously treat all fictional items as negative items. However, some of these items may exhibit semantic similarities to positive items (e.g., ”Iron Man: AI Rebellion”). Additionally, certain items might are fictional simply because they have not yet been released (e.g., ”Iron Man 4”). These cases represent incorrect negative signals.
[…]
To fully unlock the potential of LLMs in recommendation, supervised fine-tuning (SFT) is commonly applied for LLM-based recommenders (Bao et al., 2023b, a; Liao et al., 2024; Kim et al., 2024; Gao et al., 2024; Zheng et al., 2024). These methods typically structure users’ historical interactions as prompts, paired with descriptions of positive items as target responses, and fine-tune LLMs using a Language Modeling Loss (LML) (Radford, 2018). This loss, inherited from language generation tasks and expressed as a token-wise softmax loss, augments the probability (i.e., logits) of tokens representing positive items while penalizing the logits of other generated content. However, we argue that this objective has significant limitations in the recommendation scenario:
  • Significant Divergence from the Recommendation Objective: RS aims for personalized ranking performance (e.g., higher NDCG), prioritizing positive items over negative ones. LML deviates significantly from this ranking objective. Through extensive theoretical and empirical analyses, we find that optimizing LML primarily focuses on generating valid item descriptions that exist in the system, while providing limited guidance to help LLMs differentiate positive items from negative ones. This deviation significantly hinders the effectiveness of LML in recommendations.
  • Improper Negative Signals: Language modeling loss implicitly considers all other generated item descriptions as negative, including valid negative items that the user has not interacted with and fictitious items that do not exist in the RS. This treatment is flawed as it is improper to hypothesize that the user dislikes these fictitious items. In fact, some fictitious items may share similar semantic with the positive items, whose contents may be favored by users. As shown in Figure 1, a typical fan of the Marvel Universe who has watched ”Iron Man 1” would likely enjoy a fictitious movie such as ”Iron Man: AI Rebellion”. Blindly treating all such fictitious items as negative could confuse the LLM in capturing user preferences.
To tackle these limitations, we introduce a novel loss function, termed Masked Softmax Loss (MSL), specifically designed for fine-tuning LLMs as recommenders. MSL employs a masking mechanism that prevents penalization of fictitious item descriptions. As illustrated in Figure 2, this mechanism can be implemented efficiently at the token level by masking the invalid tokens in the softmax calculation that correspond fictitious items. Our theoretical analyses further demonstrate the close connection of MSL with the ranking objective, serving as a tight upper bound of the NDCG metrics.
Despite its theoretical advantages, MSL may encounter gradient vanishing issues during practical application. This arises from the reduced number of tokens in the softmax denominator, which can lead to particularly small gradients and loss values. A simple and effective strategy to address this is the introduction of an additional hyperparameter, temperature, in the softmax function to modulate its values. While effective, this approach requires tedious and time-consuming hyperparameter tuning, which is unsatisfactory for LLM-based recommendations. To address this challenge, we propose an Adaptive Temperature Strategy (ATS). By examining the gradient of MSL and the role of temperature, we derive an adaptive configuration based on the average number of valid tokens in the dataset. This strategy effectively mitigates gradient vanishing in MSL without requiring extensive hyperparameter tuning.
Figure 2. The schematic diagram illustrates Language Modeling Loss (LML) and Masked Softmax Loss (MSL).
[…]
Lastly, in terms of aligning LLM with the ranking objective, the most relevant work is the recently proposed S-DPO (Chen et al., 2024b), which integrates Direct Preference Optimization (DPO) (Rafailov et al., 2024) in LLM-based recommendation. However, S-DPO exhibits several limitations: 1) Suboptimal Performance: S-DPO still relies on LML to fine-tune the model and considers the fine-tuned LLM as a reference model for optimization. Given the inherent limitations of LML, the effectiveness of S-DPO is compromised. 2) Unstable Results: S-DPO requires negative item sampling, which can lead to training instability, especially in fine-tuning tasks with limited epochs. 3) High Computational Cost: S-DPO requires more training instances and epochs, resulting in significantly longer training times (approximately 4 times) compared to MSL.
In summary, this work makes the following contributions:
  • We propose a novel loss function, Masked Softmax Loss (MSL), specifically tailored for fine-tuning large language models to effectively align with recommendation objectives.
  • We address the potential gradient vanishing issue of MSL by developing an adaptive temperature strategy that mitigates this issue without requiring hyperparameter tuning.
  • Extensive experiments on four real-world datasets demonstrate that the proposed MSL outperforms LML by a large margin (42.24% on average in NDCG@10).
Table 1. Prompt templates for implementing recommendation tasks (using Toy dataset as an example)
Instruction Input
Instruction: Given a list of toys the user has played before, please recommend a new toy that the user likes to the user.
Input: The user has played the following toys before: ”LeapFrog Discovery Ball”, ”Plush Elmo Knows Your Name”, ”Blokus Game”, …
Instruction Output
Output: ”MindWare Q-Ba-Maze Cool Colors”
2. LLM-based Recommendation
Referring to recent work (Liao et al., 2024; Bao et al., 2023a; Na et al., 2024; Bao et al., 2023b; Lin et al., 2024b; Zheng et al., 2024), this work also focuses on sequential recommendation, which holds notable practical significance by considering the temporal order of user behavior. Given a sequential recommender system with a user set $\mathcal{U}$ and an item set $\mathcal{V}$, let user’s historical interactions be denoted as $S=\{s_{1},s_{2},...\}$, where $s_{i}\in\mathcal{V}$ denotes the $i$-th interacted item in the sequence. The objective of sequential recommendation is to infer user preferences from $S$ and retrieve the positive item $p$ that the user will interact with next. This task is often conceptualized as a ranking problem, aiming to position the positive item $p$ higher in the ranking list. Consequently, ranking metrics such as NDCG are frequently adopted to evaluate recommendation performance.
Given the remarkable success of large language models (LLMs) across various domains (Wu et al., 2023b; Ouyang et al., 2022; Tang et al., 2024), integrating LLMs into recommendation systems has been extensively explored (Wu et al., 2024). A prominent strategy is to directly leverage powerful LLMs as recommenders. As shown in Table 1, this paradigm organizes users’ historical interactions as language prompts $x$, typically consisting of the descriptions (e.g., titles) of the items in $S$ and the description of the recommendation task. This prompt is then used to instruct the LLMs to predict the item (descriptions) that the user is most likely to interact with.
Since LLMs are typically not pre-trained on recommendation data, supervised fine-tuning is necessary to align LLMs with the recommendation task. This strategy pairs the prompts $x$ and the description of the target positive item $y^{p}$ as a training instance $(x,y^{p})$, and optimizes LLMs with the following Language Modeling Loss (LML):
(1) $\mathcal{L}_{LML}(x,y^{p};\theta)$ $=-\log P_{\theta}\left(y^{p}\mid x\right)=\sum_{t=1}^{|y^{p}|}-\log P_{\theta}\left(y^{p}_{t}\mid x,y^{p}_{<t}\right)$
$=\sum_{t=1}^{|y^{p}|}-\log\frac{\exp(f_{\theta}(y_{t}^{p}|x,y^{p}_{<t}))}{\sum_{z\in\mathcal{Z}}\exp(f_{\theta}(z|x,y^{p}_{<t}))}$
where $y^{p}_{t}$ denotes the $t$-th token of the positive item description $y^{p}$, and $y^{p}_{<t}$ represents the token sequence preceding $y^{p}_{t}$. The set $\mathcal{Z}$ corresponds to the entire vocabulary of tokens in the LLM, and $f_{\theta}(y^{p}_{t}\mid x,y^{p}_{<t})$ denotes the logit of the token $y^{p}_{t}$ predicted by LLMs, where $\theta$ denotes the parameters of LLMs. For simplicity, we use $f_{\theta}(y^{p}_{t})$ (or $f_{\theta}(z)$) to represent $f_{\theta}(y^{p}_{t}\mid x,y^{p}_{<t})$ (or $f_{\theta}(z\mid x,y^{p}_{<t})$), and $P_{\theta}(y^{p}_{t})$ to denote $P_{\theta}(y^{p}_{t}\mid x,y^{p}_{<t})$.
The language modeling loss is directly inherited from language generation tasks, aiming to maximize the probability of the descriptions of positive items over the whole generative content space. It can be expressed in a token-wise manner with softmax loss, which augments the logits of the tokens representing positive items (numerator), while decreasing the logits of the other tokens in the vocabulary (denominator).
The notation table is presented in Table LABEL:notation.
Table 2. Notations in the paper.
Notations Descriptions
$\mathcal{U}$ user set
$\mathcal{V}$ item set
$S$ the user historical interaction sequence
$p$ the positive item of the sequence $S$
$x$ the input prompt of the sequence $S$
$y^{v}$ the description of the item $v$
$y^{v}_{t}$ $t$-th token of $y^{v}$
$\mathcal{Z}$ the vocabulary of LLM
$\mathcal{Z}_{valid}(y^{v}_{<t})$ valid tokens for a given prefix $y^{v}_{<t}$
$\theta$ the model parameter
$f_{\theta}$ logits output by the model
$P_{\theta}(y^{v}_{t})$ the probability of $y^{v}_{t}$ over $\mathcal{Z}$
$P_{\theta}^{valid}(y^{v}_{t})$ the probability of $y^{v}_{t}$ over $\mathcal{Z}_{valid}(y^{v}_{<t})$
$\mathcal{L}_{LML}$ language modeling loss
$\mathcal{L}_{MSL}$ masked softmax loss
3. Analyses on Language Modeling Loss
While language modeling loss is commonly used for fine-tuning LLMs as recommenders, we argue that it still suffers from the following limitations:
Limitation 1: Significant Divergence from the Recommendation Objective. Recommender systems aim to retrieve positive items from the valid item set in the system. In contrast, LML aims to retrieve positive item descriptions from the entire language space that LLMs could generate. It’s important to note that the language space contains descriptions of both valid items in the system and fictitious items imagined by LLMs. This causes the objective of LML to deviate significantly from the recommendation objective.
To better understand this deviation, we can decompose LML into two components:
(2) $\mathcal{L}_{LML}(x,y^{p};\theta)$ $=\mathcal{L}_{LML}^{1}(x,y^{p};\theta)+\mathcal{L}_{LML}^{2}(x,y^{p};\theta)$
(3) $\mathcal{L}_{LML}^{1}(x,y^{p};\theta)$ $=\sum_{t=1}^{|y^{p}|}\underbrace{-\log\frac{\sum_{z\in\mathcal{Z}_{valid}(y^{p}_{<t})}\exp(f_{\theta}(z)))}{\sum_{z\in\mathcal{Z}}\exp(f_{\theta}(v)))}}_{\text{Lifting valid items over invalid items}}$
(4) $\mathcal{L}_{LML}^{2}(x,y^{p};\theta)$ $=\sum_{t=1}^{|y^{p}|}\underbrace{-\log\frac{\exp(f_{\theta}(y_{t}^{p})))}{\sum_{z\in\mathcal{Z}_{valid}(y^{p}_{<t})}\exp(f_{\theta}(z)))}}_{\text{Lifting the positive items over negative items}}$
where $z\in\mathcal{Z}_{valid}(y^{p}_{<t})$ denotes a valid token, ensuring that the combined language contents $[y^{p}_{<t},z]$ can be a prefix of any valid item description. A similar definition applies to invalid tokens, $z\notin\mathcal{Z}_{valid}(y^{p}_{<t})$, which would make the generated contents fall outside the scope of valid items’ descriptions. For simplicity, we use $\mathcal{Z}_{valid}$ to denote $\mathcal{Z}_{valid}(y^{p}_{<t})$ in the following text.
Figure 3. The gradient norms of the two components of language modeling loss during the model training.
[…]
Figure 4. The variation of $\mathcal{L}_{LML}^{2}$ during the training when optimizing with $\mathcal{L}_{LML}$ vs. $\mathcal{L}_{LML}^{2}$.
[…]
Language modeling loss has two-fold effects: 1) $\mathcal{L}_{LML}^{1}$ increases the logits of valid tokens while penalizing the invalid tokens. This component would lift the probability of valid items over fictitious items, guiding the LLMs towards outputting a valid item description. 2) $\mathcal{L}_{LML}^{2}$ increases the logits of the positive token (i.e., $y^{p}_{t}$) and penalizes the logits of negative tokens (i.e., $\mathcal{Z}_{valid}\setminus y^{p}_{t}$). This component would lift the probability of positive items over negative items, which aligns with the recommendation objective. We will also prove the close theoretical connection of this component with the NDCG metrics in the next section (lemma 4.1).
The above decomposition illustrates the differences and connections between LML and the recommendation objective. While the recommendation objective serves as one component of LML, we empirically find that optimizing LML is ineffective, as the gradient is dominated by $\mathcal{L}_{LML}^{1}$. Figure 3 illustrates this point, showing the norm of the gradient from two components on typical datasets Toy and Book. It can be observed that $\mathcal{L}_{LML}^{1}$ exerts an overwhelming effect on the training, hindering the convergence of $\mathcal{L}_{LML}^{2}$. To further demonstrate this point, we conduct another experiment as shown in Figure 4, where we visualize the training loss of $\mathcal{L}^{2}_{LML}$ when we optimize $\mathcal{L}_{LML}$ or $\mathcal{L}_{LML}^{2}$ only for comparison. As can be seen, the training loss when we directly optimize $\mathcal{L}_{LML}^{2}$ decreases quickly, while the loss drop under optimizing $\mathcal{L}_{LML}$ seems hindered. These analyses demonstrate the ineffectiveness of leveraging LML in improving recommendation performance.
Limitation 2: Improper Negative Signals. From Eq.(2), we find that LML penalizes the logits of invalid tokens, implicitly considering all fictitious items that are not exist on the system as negative items. However, this treatment is flawed as it is improper to hypothesize that the user dislikes these fictitious items. In fact, some fictitious items may share semantic similarities with positive items and could potentially align with user preferences. To illustrate this issue, consider the example of a typical fan of the Marvel Universe who enjoys the movie ”Iron Man 1” as shown in Figure 1. Some of the fictitious items may share semantic similarities with the positive items and may be favored by users (e.g., ”Iron Man: AI Rebellion”). Additionally, certain items are fictional simply because they have not yet been released (e.g., ”Iron Man 4”). As such, blindly treating all such fictitious items as negative could confuse the LLM, giving incorrect signals for capturing user preference.
Table 3. The deviation of NDCG@10 under multiple random seeds.
Method Toy Book
S-DPO $0.0219\pm 0.0036(16.4\%)$ $0.0124\pm 0.0019(15.3\%)$
MSL $0.0294\pm 0.0006(2.0\%)$ $0.0175\pm 0.0005(2.9\%)$
Analyses on S-DPO. While S-DPO (Chen et al., 2024b) leverages direct preference optimization to enhance LLM-based recommendation, it still suffers from the following limitations:
  • Suboptimal Performance. S-DPO can not address the aforementioned limitations inherent in LML. S-DPO still relies LML to fine-tune LLMs, which would be utilized as a reference model for further DPO optimization. Given the inherent limitations of LML, the effectiveness of S-DPO is compromised (cf. Section 5.2).
  • Unstable Performance. S-DPO relies on sampling negative items to establish the ranking relationship between positive and negative items. However, this sampling process incurs performance instability. To evaluate this, we train the model using multiple random seeds and calculated the deviation of NDCG@10. As shown in Table 3, S-DPO exhibits significantly higher deviation compared to our proposed MSL, with performance losses reaching up to 16.4% on the Toy dataset and 15.3% on the Book dataset. This highlights the instability of S-DPO’s performance.
  • High Computational Cost. S-DPO requires further fine-tuning on the reference model, which entails additional training epochs. Furthermore, the inclusion of extra negative items substantially increases the data size. These factors contribute to its inefficiency. Empirically, S-DPO requires nearly four times the runtime of MSL and LML (cf. Section 5.4).
4. Methodology
In this section, we first detail the proposed Masked Softmax Loss (MSL) to address the limitations of language modeling loss (Subsection 4.1). We then highlight the potential gradient vanishing challenge in MSL and propose the Adaptive Temperature Strategy to tackle this issue (Subsection 4.2). The schematic diagrams of the MSL and LML methods are shown in Figure 2.
4.1. Masked Softmax Loss
The above analyses reveal that the limitations of LML primarily lie in the penalization of invalid tokens — it not only causes the loss to deviate from the recommendation objective but also introduces improper negative signals. To address this, a straightforward approach is to mask the invalid tokens in LML, i.e., directly leverage the second component of LML to optimize LLMs. Formally, the Masked Softmax Loss is formulated as follows:
(5) $\mathcal{L}_{MSL}(x,y^{p};\theta)=\sum_{t=1}^{|y^{p}|}-\log\frac{\exp(f_{\theta}(y_{t}^{p})))}{\sum_{z\in\mathcal{Z}_{valid}}\exp(f_{\theta}(z)))}$
This simple strategy effectively addresses the limitations by eliminating the penalization of invalid tokens. One might be concerned that this strategy could increase the risk of hallucination (Ji et al., 2023), where LLMs generate fictitious item descriptions during the inference stage. This concern can be easily mitigated by employing constrained beam search during generation (De Cao et al., 2020). Specifically, when choosing or sampling the next token in beam search, the selection can be restricted to valid tokens rather than the entire vocabulary. Such strategy ensures that the generated content corresponds to a valid item in the system, effectively mitigating the hallucination issue.
Overall, MSL possesses the following desirable properties:
Alignment with the Recommendation Objective. Intuitively, masking invalid tokens guide the model to focus more on differentiating positive items from negative ones. In fact, we have the following lemma establishing the theoretical connections between MSL and NDCG:
Lemma 0.
Considering a LLM-based RS that leverages the scores $P_{\theta}^{valid}(y^{v}|x)=\prod_{t=1}^{|y^{v}|}P_{\theta}^{valid}(y^{v}_{t}|x,y^{v}_{<t})$ for ranking items, where
$P_{\theta}^{valid}(y^{v}_{t}|x,y^{v}_{<t})=\frac{\exp(f_{\theta}(y_{t}^{v})))}{\sum_{z\in\mathcal{Z}_{valid}(y^{v}_{<t})}\exp{(f_{\theta}(z))}}$
represents the probability of the token $y^{v}_{t}$ within the valid token set $\mathcal{Z}_{valid}(y^{v}_{<t})$, optimizing $\mathcal{L}_{MSL}(x,y^{p};\theta)$ serves as a tighter upper bound of $-logNDCG(S)$ compared with $\mathcal{L}_{LML}(x,y^{p};\theta)$, i.e.,
$-logNDCG(S)\leq\mathcal{L}_{MSL}(x,y^{p};\theta)\leq\mathcal{L}_{LML}(x,y^{p};\theta)$
The proof is presented in appendix A. Note that the premise of ranking items based on $P_{\theta}^{valid}(y^{v}|x)$ is naturally satisfied when we mask invalid tokens during generation using constrained beam search with a large beam size. This lemma demonstrates that MSL is well-aligned with the recommendation objective and provides a tighter upper bound for optimizing NDCG compared to LML. $\mathcal{L}_{LML}^{1}$ in LML is redundant for NDCG optimization and may even introduce interference as previously discussed. Consequently, MSL is theoretically anticipated to achieve superior performance.
Ease of Implementation. Our MSL is simple, easily implemented, and can serve as a suitable surrogate for LML with minimal code revisions. The main implementation complexity lies in identifying valid tokens. In fact, this can be easily achieved by using a trie tree (a.k.a. a prefix tree) (Bodon and Rónyai, 2003). We can utilize existing packages of marisa-trie (with only 3 lines of codes) to construct the trie tree from all item descriptions during the pre-processing stage and calculate the masking matrix. Subsequently, we can revise the LML to MSL by simply applying the masking matrix. MSL can be seamlessly integrated into various existing LLM-based recommendation methods, including the recently proposed BIGRec (Bao et al., 2023a), LLaRA (Liao et al., 2024), A-LLM (Kim et al., 2024), and consistently yield improvements (cf. section 5.2).
Efficiency. The primary computational challenge lies in constructing the trie tree and the masking matrix. However, this process is highly efficient, with a time and memory complexity of $O(|\mathcal{V}|\hat{l})$, where $|\mathcal{V}|$ denotes the number of items in the system, and $\hat{l}$ represents the average token length of the item description. Empirically, the Trie tree construction for all datasets is completed in under one second. Furthermore, MSL improves efficiency by excluding invalid tokens from the loss calculation (cf. section 5.4).
4.2. Adaptive Temperature Strategy
4.2.1. Potential Gradient Vanishing Issue
Despite the theoretical advantages of MSL, it may encounter gradient vanishing issues in practical applications, challenging its effectiveness. To illustrate this effect, the gradient of MSL over each sample can be expressed as follows:
(6) $\nabla_{\theta}\mathcal{L}_{MSL}(x,y^{p};\theta)$ $=-\sum_{t=1}^{|y^{p}|}w(y_{t}^{p})g(y_{t}^{p},\theta)$
where
(7) $w(y_{t}^{p})=1-P^{valid}_{\theta}(y_{t}^{p}\mid x,y^{p}_{<t})$
(8) $g(y_{t}^{p},\theta)=\nabla_{\theta}f_{\theta}(y_{t}^{p})-\frac{\sum_{z\in\mathcal{Z}^{\prime}}\exp(f_{\theta}(z))\nabla_{\theta}f_{\theta}(z)}{\sum_{z\in\mathcal{Z}^{\prime}}\exp(f_{\theta}(z))}$
$\mathcal{Z}^{\prime}=\mathcal{Z}_{valid}\setminus\{y_{t}^{p}\}$ represents the set of negative tokens. For simplicity, let $P_{\theta}^{valid}(y^{p}_{t})$ represent $P_{\theta}^{valid}(y^{p}_{t}\mid x,y^{p}_{<t})$.
Figure 5. The weight value $w(y_{t}^{p})$ distribution of a batch of samples for MSL and LML. The weight are sorted in descending order.
[…]
As observed, the magnitude of gradient is influenced by the weight $w(y_{t}^{p})$. The gradient vanishing phenomenon can be attributed to a reduced number of terms in the denominator of $P_{\theta}^{valid}(y_{t}^{p})$ as compared with $P_{\theta}(y_{t}^{p})$, which naturally increases $P_{\theta}^{valid}(y_{t}^{p})$, thereby decreasing the weight $w(y_{t}^{p})$. This reduction can even cause the gradient to approach zero, particularly because the logits of positive tokens are often larger than those of other valid tokens [^1] This assumption is reasonable, as the optimization process tends to increase the logits of positive tokens while decreasing those of negative tokens..
Figure 5 presents the weight values of a batch of samples for both LML and MSL. When the valid token mask is applied to MSL, the weight values of all samples are substantially reduced, with some values even nearing zero. Importantly, these tokens are often crucial, as they are typically located among the first few tokens in the response and play a pivotal role in training. Empirical analysis on the Office dataset underscores this point: 61% of samples with weight values below 0.1 are concentrated within the first three tokens of the item. Similar patterns are observed across other datasets.
4.2.2. The Introduction of Temperature
To tackle this issue, we have found that the introduction of a temperature $\tau$ can effectively address this problem:
(9) $\mathcal{L}_{MSL}(x,y^{p};\theta)=\sum_{t=1}^{|y^{p}|}-\tau\log\frac{\exp(f_{\theta}(y_{t}^{p})/\tau))}{\sum_{z\in\mathcal{Z}_{valid}}\exp(f_{\theta}(z)/\tau))}$
where the weight $w(y_{t}^{p})$ can be written as:
(10) $w(y_{t}^{p})=1-P^{valid}_{\theta}(y_{t}^{p})=1-\frac{\exp(f_{\theta}(y_{t}^{p})/\tau))}{\sum_{z\in\mathcal{Z}_{valid}}\exp{(f_{\theta}(z)/\tau)}}$
The introduction of temperature can modulate the magnitude of the gradient. Considering that the logits of positive tokens are typically larger than those of other tokens, an increase in $\tau$ would relatively reduce the value of $P^{valid}_{\theta}(y_{t}^{p})$, increasing $w(y_{t}^{p})$; Figure 6 highlights the impact of incorporating temperature, which leads to a significant improvement in performance. Conversely, alternative approaches, such as adjusting the learning rate or introducing a balancing coefficient for negative tokens, fail to yield satisfactory results. The empirical evidence supporting these findings will be presented in Section 5.3.
Figure 6. The performance of MSL across different temperature $\tau$.
[…]
4.2.3. Adaptive Temperature Strategy
Despite its effectiveness, the introduction of $\tau$ incurs another hyperparameter tuning challenge. Given that the average number of valid tokens varies across different datasets, the optimal value of $\tau$ naturally evolves. For example, the optimal $\tau$ on dataset Book is 1.5, while it is 4.5 on Toy as shown in Figure 6. Transferring the optimal $\tau$ from one dataset to another without adjustment can lead to significant performance drops. This necessitates extensive hyperparameter tuning of $\tau$, which can be particularly time-consuming, especially for heavy LLM-based recommenders.
To address this, inspired by recent studies on temperature (Chen et al., 2023), we develop an Adaptive Temperature Strategy (ATS) for MSL. This strategy dynamically and adaptively adjusts $\tau$ to ensure that $P_{\theta}^{valid}(y_{t}^{p})$ remains close to a target value $\eta$, preventing it from becoming excessively large and incurring gradient vanishing. Specifically, we have the following lemma:
Lemma 0.
For each training token instance $(x,y_{t}^{p})$, assuming the logits of the valid tokens $f_{\theta}(z),z\in\mathcal{Z}_{valid}$ follow a Gaussian distribution $\mathcal{N}(\mu_{t},\sigma_{t}^{2})$. Then $\tau_{t}$ for the equation $P_{\theta}^{valid}(y_{t}^{p})=\eta$ can be approximated as:
(11) $\tau_{t}\approx\frac{(f_{\theta}(y_{t}^{p})-\mu_{t})-\sqrt{(f_{\theta}(y_{t}^{p})-\mu_{t})^{2}-2\sigma^{2}_{t}\log(|\mathcal{Z}_{valid}|\eta)}}{2\log(|\mathcal{Z}_{valid}|\eta)}$
The proof is presented in the appendix B. The proof references the work (Chen et al., 2023) but adapts the process to token-wise LLM-based recommendation scenarios and different distribution conditions. The assumption of a Gaussian distribution nearly holds, as discussed in the appendix C.
Eq.(11) gives the token-wise optimal configuration of $\tau_{t}$. To make the training more stable and reduce the extra effort of calculating the token-wise $\tau_{t}$, we prefer to set a global uniform $\tau$ across various training instances:
(12) $\tau\approx\frac{(f_{\theta}(y_{t}^{p})-\mu)-\sqrt{(f_{\theta}(y_{t}^{p})-\mu)^{2}-2\sigma^{2}\log m\eta}}{2\log m\eta}$
where $\mu$, $\sigma^{2}$ denote the mean and variance of $f_{\theta}(z)$ for all training instances, and $m$ denotes the average number of valid tokens. This Eq.(12) can adaptively adjust the value of $\tau$ according to the current model state and the average number of valid tokens in the datasets, serving as an efficient alternative to brute-force hyperparameter search.
Table 4. Statistics of the datasets. AVT represents the average number of valid tokens per token instance.
Dataset #Users #Items #Interactions #Density #AVT
Toy 19124 11758 165247 0.0735% 54.38
Book 16559 6344 151928 0.1446% 70.90
Clothing 39230 22948 277534 0.0308% 53.74
Office 4895 2414 53149 0.4498% 8.98
5. Experiments
We aim to answer the following research questions:
  • RQ1: How does MSL perform compare to existing state-of-the-art recommendation methods?
  • RQ2: How do different components of MSL affect?
  • RQ3: How does MSL perform compared with state-of-the-art in terms of both accuracy and efficiency?
Table 5. The performance comparison on four real-world datasets. The best result is bolded. Improvement denotes the improvement of MSL over the best results obtained using LML and S-DPO. ”N” represents NDCG, and ”H” represents Hit Ratio.
Toy Book Clothing Office
Method N@5 N@10 H@5 H@10 N@5 N@10 H@5 H@10 N@5 N@10 H@5 H@10 N@5 N@10 H@5 H@10
SASRec 0.0101 0.0126 0.0190 0.0265 0.0097 0.0133 0.0176 0.0285 0.0046 0.0056 0.0086 0.0116 0.0132 0.0183 0.0260 0.0421
BERT4Rec 0.0157 0.0191 0.0229 0.0336 0.0118 0.0171 0.0187 0.0351 0.0071 0.0093 0.0110 0.0180 0.0225 0.0307 0.0358 0.0618
SASRec+DROS 0.0129 0.0160 0.0217 0.0311 0.0110 0.0156 0.0196 0.0340 0.0050 0.0067 0.0088 0.0142 0.0130 0.0226 0.0260 0.0561
LLM-CF 0.0103 0.0132 0.0186 0.0275 0.0106 0.0142 0.0178 0.0292 0.0041 0.0052 0.0074 0.0116 0.0144 0.0192 0.0239 0.0395
DLLM4Rec 0.0104 0.0134 0.0190 0.0284 0.0099 0.0136 0.0178 0.0303 0.0042 0.0061 0.0082 0.0138 0.0137 0.0198 0.0239 0.0504
BIGRec+LML 0.0138 0.0182 0.0213 0.0353 0.0109 0.0137 0.0169 0.0258 0.0047 0.0073 0.0092 0.0174 0.0113 0.0220 0.0203 0.0530
BIGRec+SDPO 0.0174 0.0219 0.0271 0.0413 0.0118 0.0145 0.0189 0.0276 0.0062 0.0089 0.0114 0.0198 0.0129 0.0256 0.0239 0.0629
BIGRec+MSL 0.0245 0.0288 0.0357 0.0488 0.0125 0.0172 0.0214 0.0356 0.0091 0.0120 0.0146 0.0236 0.0402 0.0438 0.0556 0.0665
Improvement 41.06% 31.35% 31.54% 18.18% 5.70% 18.55% 13.25% 28.93% 46.05% 34.52% 28.07% 19.19% 213.17% 70.80% 132.61% 5.79%
LLaRA+LML 0.0145 0.0193 0.0225 0.0375 0.0102 0.0132 0.0173 0.0265 0.0050 0.0083 0.0088 0.0190 0.0093 0.0204 0.0177 0.0514
LLaRA+SDPO 0.0164 0.0211 0.0250 0.0394 0.0094 0.0114 0.0128 0.0187 0.0077 0.0099 0.0138 0.0206 0.0196 0.0242 0.0379 0.0524
LLaRA+MSL 0.0233 0.0283 0.0336 0.0488 0.0135 0.0175 0.0244 0.0365 0.0108 0.0140 0.0176 0.0274 0.0374 0.0417 0.0535 0.0670
Improvement 41.98% 34.05% 34.17% 23.81% 31.96% 32.71% 40.79% 37.93% 40.12% 41.61% 27.54% 33.01% 90.87% 72.08% 41.10% 27.82%
A-LLM+LML 0.0151 0.0197 0.0236 0.0378 0.0107 0.0132 0.0173 0.0251 0.0055 0.0082 0.0100 0.0182 0.0110 0.0219 0.0192 0.0519
A-LLM+SDPO 0.0155 0.0201 0.0248 0.0388 0.0065 0.0093 0.0094 0.0180 0.0072 0.0104 0.0116 0.0216 0.0189 0.0240 0.0306 0.0468
A-LLM+MSL 0.0248 0.0296 0.0365 0.0513 0.0130 0.0184 0.0228 0.0395 0.0097 0.0127 0.0158 0.0252 0.0353 0.0388 0.0488 0.0597
Improvement 59.81% 47.73% 47.06% 32.26% 21.01% 39.44% 31.58% 57.27% 34.65% 22.54% 36.21% 16.67% 86.84% 61.46% 59.32% 15.00%
5.1. Experimental Settings
5.1.1. Datasets
Four conventional real-world datasets: Amazon Toys and Games, Amazon Books, Amazon Clothing, Shoes and Jewelry and Amazon Office Products  [^2] https://jmcauley.ucsd.edu/data/amazon/index_2014.html are utilized in our experiments, which are commonly used for the studies of LLM-based recommendation (Cui et al., 2024; Bao et al., 2023a; Cao et al., 2024; Lee et al., 2024; Li et al., 2023a). To ensure a fair comparison, we adopt the same data preprocessing used in recent studies (Bao et al., 2023a; Cui et al., 2024). Specifically, we firstly apply the 5-core setting to the original dataset, then for user interaction sequences longer than 11 interactions, a sliding window of length 11 is applied to segment the sequences. The resulting sequences are then sorted in ascending order by timestamp and split into training, validation, and testing sets with an 8:1:1 ratio. We randomly retain 100,000 items for Amazon Books before 5-core processing due to its large size. The processed dataset statistics are presented in Table 4.
5.1.2. Baselines
The methods compared fall into several categories:
  • Traditional recommenders (SASRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019), DROS (Yang et al., 2023)) SASRec utilizes a self-attention-based model to capture user interests. BERTRec adopts the BERT to bidirectional model user preferences. DROS incorporates DRO to improve the model’s resilience to distributional shifts.
  • LLM-enhanced recommenders (DLLM2Rec (Cui et al., 2024), LLM-CF (Sun et al., 2024)) DLLM2Rec introduces a distillation module designed to bridge the performance gap between LLMs and traditional RS. LLM-CF enhances traditional RS by integrating reasoning-driven collaborative filtering features derived from LLMs using CoT techniques. We use SASRec as the backbone for LLM-enhanced recommenders.
  • LLM-based recommenders (BIGRec (Bao et al., 2023a), LLaRA (Liao et al., 2024), A-LLM (Kim et al., 2024)) BIGRec develops instruction-tuning templates to fine-tuning LLMs on RS datasets. LLaRA enhances collaborative signals by incorporating embeddings produced by traditional models into prompts. A-LLM further alignes these embeddings with corresponding textual information.
  • Improved Loss Function for LLM-based Recommenders (S-DPO (Chen et al., 2024b)) S-DPO leverages the DPO to guide LLMs using the ranking information of positive and negative samples.
5.1.3. Implementation Details
LLaMA3 8B model (Dubey et al., 2024) is utilized as the backbone of all the LLM-based recommenders. As for training LLM-based recommenders, we train the models for 10 epochs and report the results of the epoch with the highest NDCG@5. For inference, we evaluate two mainstream methods as baselines: grounding (Bao et al., 2023a) and constrained beam search (Zhu et al., 2024), and we report the better-performing results. The ranking results obtained from constrained beam search are used to construct the recommendation list, with the number of beams fixed at 10. For MSL, we only modify the loss function of the original backbone while following its original hyperparameter settings. The parameter $\eta$ is set to 0.25. To ensure fair comparisons, we leverage the source code provided in the original papers and tune the hyperparameters of all baseline methods following the guidelines specified in their respective works. Two widely-used metrics NDCG@K and Hit Ratio@K are employed for evaluating the recommendation accuracy (K = 5, 10).
5.2. Performance Comparison (RQ1)
Table 5 provides a comparative analysis of the performance of the proposed MSL method against baseline approaches.
MSL demonstrates a significant enhancement in the performance of various LLM-based recommenders. MSL consistently outperforms all baseline across all datasets. This remarkable improvement can be attributed to the design of MSL as a specialized loss function tailored for LLM-based recommenders.
The performance improvements of LLM-enhanced recommenders remain relatively limited. LLM-CF exhibits negative gains on three out of the four datasets. This underperformance is primarily due to the inherent gap between LLMs and traditional models, which hinders the effective transfer of knowledge. DLLM2Rec, which directly generates ranking results using LLMs and incorporates a specially designed distillation mechanism, partially addresses this gap. However, its performance remains constrained by the limitations of LLMs’ recommendation capabilities as teacher models.
S-DPO demonstrates limited and inconsistent performance improvements in LLM-based recommendation systems. Its dependence on the reference model, coupled with instability introduced by sampling, significantly constrains its effectiveness. As a result, S-DPO often fails to deliver consistent improvements, and in some cases, even demonstrates negative performance gains. For example, NDCG@10 of A-LLM+S-DPO decreases from 0.0132 to 0.0093 on the Book dataset.
5.3. Ablation Study (RQ2)
Table 6 presents the results of ablation study. Specifically, we investigate the effects of MSL, temperature $\tau$ and the ATS module, as well as alternative approaches to mitigating the vanishing gradient problem.
Adjusting the temperature $\tau$ significantly enhances MSL performance. MSL without temperature underperforms compared to LML on certain datasets (e.g., Toy and Office) due to the vanishing gradient issue. Employing an hyperparameter search strategy for the $\tau$ effectively addresses this issue, resulting in substantial performance improvements. This highlights the critical role of $\tau$ in unlocking MSL’s potential. Notably, a similar hyperparameter search for LML revealed only marginal performance gains, suggesting that the observed improvements are attributable to MSL itself rather than the tuning of $\tau$.
The ATS module rivals or surpasses brute-force hyperparameter search in effectiveness. This demonstrates the efficacy of ATS as a dynamic optimization strategy. ATS offers better flexibility than fixed temperature and eliminates the need for exhaustive manual tuning.
Alternative strategies for mitigating vanishing gradients show limited effectiveness. We also investigate two alternative strategies to address the vanishing gradient problem in MSL: adjusting the learning rate (i.e., MSL + tuning lr) and introducing a balancing coefficient $\alpha$ for negative tokens (i.e., MSL + $\alpha$, where the negative token component in the denominator of $\mathcal{L}_{MSL}$ is multiplied by $\alpha$). $\alpha$ is set as $|\mathcal{Z}|/|\mathcal{Z}_{\text{valid}}|$. While these adjustments yielded minor improvements, the results remained significantly inferior to those achieved through temperature adjustment.
Table 6. Ablation study. Results are reported in NDCG@10. ”+tuning $\tau$” indicates performing a hyperparameter search for the temperature $\tau$. ”+tuning lr” indicates performing a hyperparameter search for the learning rate. ”+ $\alpha$” denotes introducing a coefficient for negative tokens. MSL (w/o $\tau$) represents MSL without the temperature. MSL (w/ ATS) represents MSL with the temperature adjusted using ATS.
Method Toy Book Clothing Office
LML 0.0182 0.0137 0.0073 0.0220
LML + tuning $\tau$ 0.0182 0.0146 0.0074 0.0248
MSL (w/o $\tau$) 0.0171 0.0151 0.0097 0.0135
MSL + tuning lr 0.0184 0.0151 0.0101 0.0177
MSL + $\alpha$ 0.0202 0.0157 0.0078 0.0221
MSL + tuning $\tau$ 0.0280 0.0182 0.0130 0.0418
MSL (w/ ATS) 0.0288 0.0172 0.0120 0.0438
5.4. Efficiency Comparison (RQ3)
In this section, we analyze and compare the efficiency and performance of various methods. As illustrated in Figure 7, MSL achieves optimal recommendation performance while simultaneously demonstrating superior computational efficiency. Specifically, MSL improves efficiency by 315% and 324% on the Toy and Book datasets, respectively, compared to S-DPO. In comparison to LML, MSL achieves efficiency gains of 4.4% and 6.7% on the same datasets.
The enhanced efficiency of MSL stems from its ability to significantly reduce the number of invalid tokens. MSL restricts the scope to a small subset of valid tokens (the average number of valid tokens for each dataset is shown in Table 4) compared to the 128,000 tokens in LLaMA3 vocabulary. This targeted approach minimizes computational overhead. The only additional overhead introduced by MSL arises from the construction of the Trie tree during the data preprocessing stage. As discussed in Section 4.1, this process is highly efficient. Table 7 shows that the consuming time for all datasets is consistently within one second. In contrast, S-DPO introduces additional computational overhead by requiring the sampling of $n$ negative samples for each positive sample, leading to approximately four times runtime compared to MSL.
Figure 7. Performance comparisons in terms of both recommendation accuracy and efficiency.
[…]
6. Related Work
6.1. Sequential Recommendation
Sequential recommendation aims to predict the next item of interest for a user based on their historical interactions. It has increasingly adopted various deep learning models over recent years. For example, GRU4Rec (Hidasi et al., 2015) leverages recurrent neural networks (RNNs), while Caser (Tang and Wang, 2018) utilizes convolutional neural networks (CNNs). More recently, models such as SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019) are built upon the self-attention mechanism (Vaswani, 2017; Chen et al., 2024a), which automatically assigns weights to each interaction to capture their relative importance. Additionally, DROS (Yang et al., 2023) incorporates distributionally robust optimization (DRO) (Rahimian and Mehrotra, 2019; Wu et al., 2023a) to improve the model’s robustness to out-of-distribution scenarios, which are common in RS (Wang et al., 2024a; Lin et al., 2024a; Chen et al., 2021; Zhao et al., 2022; Gao et al., 2023a, c). The readers may refer to the survey (Fang et al., 2020; Wang et al., 2019) for more details.
Table 7. Time of constructing trie tree.
Dataset Toy Book Clothing Office
Time 0.38s 0.17s 0.81s 0.23s
6.2. LLMs for Recommendation
Large language models (LLMs), with their powerful comprehension capabilities and extensive knowledge (Dubey et al., 2024; Achiam et al., 2023; Wang et al., 2024c), have been widely applied in RS. Two primary paradigms for using LLMs in RS as following.
LLM-based Recommenders. This paradigm attempts directly leveraging pre-trained LLMs as the backbone for recommendations using their zero-shot capabilities (Gao et al., 2023b; Hou et al., 2024; Liu et al., 2024a; Wang and Lim, 2023; Wang et al., 2024d). However, these methods often perform poorly due to the significant discrepancy between the recommendation tasks and the training objectives of LLMs (Yang et al., 2024a). To address this, subsequent research has reformulated recommendation data into prompt formats and fine-tuned LLMs to improve their performance, achieving better results (Bao et al., 2023a, b; Hu et al., 2024; Li et al., 2023b, 2024b, c, 2024a; Shi et al., 2024; Wang et al., 2024b; Zhang et al., 2023, 2024; Gao et al., 2024; Lin et al., 2024c; Xu et al., 2024a; Wang et al., 2023, 2024e; Zhu et al., 2024).
However, these studies all rely on the LML for fine-tuning, without addressing the inherent misalignment between LML and recommendation tasks. To address this issue, S-DPO (Chen et al., 2024b) builds upon DPO (Rafailov et al., 2024) by constructing positive and negative samples, explicitly incorporating ranking information into the model training process. However, it suffers from performance instability, limited effectiveness, and low efficiency. Our proposed MSL effectively addresses these issues and significantly improves the performance of LLM-based RS.
LLM-enhanced Recommenders. This paradigm primarily utilizes LLMs in auxiliary roles, such as encoders to embed the semantic information of users and items (Xi et al., 2023; Ren et al., 2024b; Wei et al., 2024; Wang et al., 2023; Lee et al., 2024; Ren et al., 2024a; Xu et al., 2024b), as an additional knowledge base (Yang et al., 2024b; Qin et al., 2024), as a reasoning tool to generate chain-of-thought (CoT) data (Sun et al., 2024; Wang et al., 2024f), or serves as teachers using distillation (Cui et al., 2024; Liu et al., 2024b). The main challenge of this paradigm lies in the significant gap between LLMs and traditional recommendation models, which hinders the effective transfer of knowledge.
7. Conclusion
In this paper, we introduce a novel loss function, MSL, specifically tailored for LLM-based RS. MSL excludes invalid tokens from participating in the loss calculation, achieving better alignment with the recommendation objectives and avoid the interference from erroneous negative signal. Despite its advantages, it can lead to gradient vanishing issues during training. To mitigate this, we introduce the temperature coefficient and propose an Adaptive Temperature Strategy, which adaptively adjusts the temperature without requiring extensive hyperparameter tuning. We validate the effectiveness of MSL through theoretical analysis and empirical experiments. Our findings demonstrate that MSL significantly improves the performance of state-of-the-art LLM-based recommendation models.
This study highlights the importance of optimizing LLMs for recommendation tasks by refining their loss functions. Future research could explore the design of specialized LLM architectures to further enhance their suitability for recommendation systems.
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (62372399, 62476244), OPPO Research Fund, and the advanced computing resources provided by the Supercomputing Center of Hangzhou City University.
Appendix A The proof of lemma 4.1
Proof.
(13) $\mathcal{L}_{MSL}(x,y^{p};\theta)$ $=\sum_{t=1}^{|y^{p}|}-\log\frac{\exp(f_{\theta}(y_{t}^{p})))}{\sum_{z\in\mathcal{Z}_{valid}}\exp(f_{\theta}(z)))}$
$=-\log\prod_{t=1}^{|y^{p}|}P_{\theta}^{valid}\left(y^{p}_{t}\mid x,y^{p}_{<t}\right)=-\log P_{\theta}^{valid}\left(y^{p}\mid x\right)$
Since the probabilities of all items sum to 1,
(14) $-\log P_{\theta}^{valid}\left(y^{p}\mid x\right)=-\log\frac{1}{\sum_{v\in\mathcal{V}}e^{\log P_{\theta}^{valid}\left(y^{v}\mid x\right)-\log P_{\theta}^{valid}\left(y^{p}\mid x\right)}}$
Using the inequality $\exp(a)\geq\mathbb{I}(a),a\in\mathbb{R}$,
(15) $\mathcal{L}_{MSL}(x,y^{p};\theta)$ $\geq-\log\frac{1}{\sum_{v\in\mathcal{V}}\mathbb{I}(\log P_{\theta}^{valid}\left(y^{p}\mid x\right)\geq\log P_{\theta}^{valid}\left(y^{v}\mid x\right))}$
$=-\log\frac{1}{\pi(p)}$
where $\pi(p)$ denotes the rank position of positive item $p$. Since $\log(1+\pi(p))\leq\pi(p)$,
(16) $-\log\frac{1}{\pi(p)}\geq-\log\frac{1}{\log(1+\pi(p))}=-\log NDCG(S)$
Since the denominator of the softmax function in each term of $\mathcal{L}_{{MSL}}$ is smaller than $\mathcal{L}_{LML}$, it follows that $\mathcal{L}_{MSL}<\mathcal{L}_{LML}$, then
(17) $-\log NDCG(S)\leq\mathcal{L}_{MSL}\leq\mathcal{L}_{LML}$
Appendix B The proof of lemma 4.2
Proof.
The denominator of $P^{valid}_{\theta}(y_{t}^{p})$, $\mathbb{E}_{f}[\exp(f/\tau_{t})]$, can be expanded as follows:
[…]
The integral corresponds to a standard Gaussian distribution is 1. Therefore,
(19) $\mathbb{E}_{f}[\exp(f/\tau_{t})]=\exp(\mu_{t}/\tau_{t})\exp(\sigma_{t}^{2}/2\tau_{t}^{2}).$
Thus, we approximate:
(20) $\mathbb{E}_{f}[\exp(f/\tau_{t})]\approx\exp(\mu_{t}/\tau_{t})\left(1+\frac{\sigma_{t}^{2}}{2\tau_{t}^{2}}\right),$
since $\exp(\sigma_{t}^{2}/2\tau_{t}^{2})=1+\frac{\sigma_{t}^{2}}{2\tau_{t}^{2}}+o((\sigma_{t}^{2}/2\tau_{t}^{2})^{2})$. Then the original equation can be approximated as:
(21) $P^{valid}_{\theta}(y_{t}^{p})\approx\frac{\exp(f_{\theta}(y_{t}^{p})/\tau_{t})}{|\mathcal{Z}_{valid}|\exp(\mu_{t}/\tau_{t}+\sigma_{t}^{2}/2\tau_{t}^{2})}=\eta.$
Solving the above quadratic equation,
(22) $\tau_{t}\approx\frac{(f_{\theta}(y_{t}^{p})-\mu_{t})-\sqrt{(f_{\theta}(y_{t}^{p})-\mu_{t})^{2}-2\sigma_{t}^{2}\log(|\mathcal{Z}_{valid}|\eta)}}{2\log(|\mathcal{Z}_{valid}|\eta)}$
Appendix C Validation of Gaussian distribution for logits
Figure 8 indicates that the logits distribution closely follows a Gaussian distribution.
Figure 8. The logits distribution and the fitting results using a Gaussian distribution on Toy and Book dataset.
[…]